OpenStudy (anonymous):

Cheyenne has z candy bags. The bag is filled with 20 gummy bears. She then gave out y amount of gummy bears to each of her friends at band camp. she has x pieces of gummy bears left. Derive an equation that shows the number of friends Cheyenne has.

OpenStudy (anonymous):

Number of friends Cheyenne has at band camp i miss that part sorry

OpenStudy (dan815):

Total gummy bears = 20*Z Let F be the number of friends she has she gave out y about to each so 20*Z-y*F= x

OpenStudy (dan815):

we want to solve for F

OpenStudy (dan815):

F=(20z-x)/y

OpenStudy (anonymous):

oh thank you i think that is same i got. can i ask you one more pleas?

OpenStudy (dan815):

oh okay next time just show the work u did so i can check

OpenStudy (anonymous):

ok this one i dont know If the distance Tommy traveled remains constant and then increases it by 25%, what was the amount of time that passed must decrease by how much percent?

OpenStudy (anonymous):

amount of time that passed decreased by what percent

OpenStudy (dan815):

okay let d be the distance tommy tavels let t be the initial time let v he the inital speed before 25% increase v=d/t t=d/v now the speed got increase bt 25% so new_v=1.25*v =5/4 v new_t=d/1.25v =d/(5v/4) =4d/5v =4/5* d/v so the new time is 4/5th of the old time 4/5 = 80% so thats a 20% decrease

OpenStudy (anonymous):

how u get 5/4?